OpenStudy (anonymous):

6x+2y=10 how to solve and get order pair?

OpenStudy (anonymous):

3x+y=5

OpenStudy (anonymous):

you can use the substitution method then when you find one varable plug it in

OpenStudy (anonymous):

1) find particular solution

OpenStudy (anonymous):

like (1,2)

OpenStudy (anonymous):

3x+y=5 3(1)+2=5 subtraction 3(x-1)+(y-2)=0

OpenStudy (anonymous):

that's mean 3(x-1)=2-y 2-y=3k y=-3k+2 x=k+1 k an integer